# Building Confidence in Deep Generative Protein Design

### v1

@@ Appendix E Evaluation of Side-Chain Homology Modeling
Appendix F Molecular Dynamics Simulations
For experimental structures, crystallographic water and unnecessary small molecules were removed. For generated structures, missing side-chains were added via homology modeling (Section 3.6). After adding hydrogen atoms using Reduce2 (Grosse-Kunstleve et al., 2002) and confirming no missing atoms, each protein was centered at the origin.
@@ Appendix F Molecular Dynamics Simulations
F.1 Molecular Dynamics Setup for Stability Assessment of Generated Structures
@@
The next step involved equilibrating the solvent and ions around the protein. We chose the leap-frog integrator for the simulations and applied the LINCS algorithm to constrain hydrogen bonds. Equilibration involved two stages. In the first stage, we performed a 500 ps NVT equilibration (250,000 steps with a 2 fs time step). Temperature control was managed using the V-rescale thermostat, with the system divided into two groups: (1) protein and (2) water + ions, both set to a target temperature of 310 K to simulate physiological conditions. In the second stage, we carried out a 500 ps NPT equilibration with pressure coupling enabled. The pressure was regulated using the C-rescale method with isotropic coupling. The target pressure was 1.0 bar, with a compressibility of $4.5\times 10^{-5}$ bar-1 and a pressure coupling time constant of 0.5 ps.
@@
F.2 Molecular Dynamics Setup for Conformational Analysis of Protein-Ligand Complexes
The receptor and ligand were saved as separate coordinate files to prepare their respective topologies. The receptor topology was prepared as in Section F.1. For the ligand, hydrogen atoms were added using OpenBabel (O’Boyle et al., 2011), and topology was generated via the CGenFF server (Vanommeslaeghe et al., 2009; Vanommeslaeghe & MacKerell, 2012; Vanommeslaeghe et al., 2012). The receptor and ligand topologies, along with force-field-compatible coordinate files, were then combined to construct the complete complex system.
The MD workflow for complexes followed the same steps in Section F.1. Complexes were placed in an octahedral simulation box, energy-minimized in vacuum, solvated in water, and neutralized with Na+ and Cl- ions to 150 mM. A second energy minimization was then performed on the solvated system.
During equilibration, positional restraints were applied to the ligand to prevent unnecessary displacement in the initial stages of the simulation. Additionally, to minimize interference from temperature fluctuations of the ligand on the overall simulation, we defined two temperature coupling groups: (1) the receptor and ligand as one group, and (2) the solvent and ions as the other. Other equilibration settings followed Section F.1.
@@ F.2 Molecular Dynamics Setup for Conformational Analysis of Protein-Ligand Complexes
Appendix G Protein-ligand Blind Docking
Similarly, crystallographic water and unwanted molecules were removed from experimental structures, and missing side-chains were added to generated structures via homology modeling. Receptor structures were prepared using AutoDock Tools (Morris et al., 2009), with polar hydrogens added, Kollman charges assigned, and any missing atoms repaired. For receptors within the same family, we prepared a shared ligand file, adding hydrogen atoms and assigning Gasteiger charges. A large grid box, typically 80 to 110 Å̊$\mathring{\text{A}}$ per side, was defined to cover the entire protein surface.
@@ Appendix G Protein-ligand Blind Docking
Appendix H Generated Structures

### v2

@@ Appendix H Evaluation of Side-Chain Homology Modeling
Appendix I Molecular Dynamics and Blind Docking Simulations
Figure 11: Stability assessment of MD simulations across proteins using various metrics. Distributions of (A) RMSD, (B) radius of gyration ($R_{g}$), (C) potential energy, and (D) secondary structure counts throughout the simulation. Interquartile ranges and whiskers show metric variation; high-quality structures have medians close to experimentally derived values with narrow ranges.
For experimentally derived structures, crystallographic water and unnecessary small molecules were removed. For generated structures, missing side-chains were added via homology modeling (Section 2.5). After adding hydrogen atoms using Reduce2 (Grosse-Kunstleve et al., 2002) and confirming no missing atoms, each protein was centered at the origin.
@@ Appendix I Molecular Dynamics and Blind Docking Simulations
I.1 Molecular Dynamics Setup for Stability Assessment of Generated Structures
@@
The next step involved equilibrating the solvent and ions around the protein. We chose the leap-frog integrator for the simulations and applied the LINCS algorithm to constrain hydrogen bonds. Equilibration involved two stages. In the first stage, we performed a 500 ps NVT [^6] Constant number of particles, volume, and temperature. equilibration (250,000 steps with a 2 fs time step). Temperature control was managed using the V-rescale thermostat, with the system divided into two groups: (1) protein and (2) water + ions, both set to a target temperature of 310 K to simulate physiological conditions. In the second stage, we carried out a 500 ps NPT  [^7] Constant number of particles, pressure, and temperature. equilibration with pressure coupling enabled. The pressure was regulated using the C-rescale method with isotropic coupling. The target pressure was 1.0 bar, with a compressibility of $4.5\times 10^{-5}$ bar-1 and a pressure coupling time constant of 0.5 ps.
@@
I.2 Molecular Dynamics Setup for Conformational Analysis of Protein-Ligand Complexes
The receptor and ligand were saved as separate coordinate files to prepare their respective topologies. The receptor topology was prepared as in Section I.1. For the ligand, hydrogen atoms were added using OpenBabel (O’Boyle et al., 2011), and topology was generated via the CGenFF server (Vanommeslaeghe et al., 2009; Vanommeslaeghe & MacKerell, 2012; Vanommeslaeghe et al., 2012). The receptor and ligand topologies, along with force-field-compatible coordinate files, were then combined to construct the complete complex system.
The MD workflow for complexes followed the same steps in Section I.1. Complexes were placed in an octahedral simulation box, energy-minimized in vacuum, solvated in water, and neutralized with Na+ and Cl- ions to 150 mM. A second energy minimization was then performed on the solvated system.
During equilibration, positional restraints were applied to the ligand to prevent unnecessary displacement in the initial stages of the simulation. Additionally, to minimize interference from temperature fluctuations of the ligand on the overall simulation, we defined two temperature coupling groups: (1) the receptor and ligand as one group, and (2) the solvent and ions as the other. Other equilibration settings followed Section I.1.
Figure 12: Comparison of experimentally derived binding modes (bold boxes; first row) with predictions from blind docking simulations for receptors binding to family-specific ligands. Ligands are colored as green (ground truth), cyan (predicted using experimentally derived receptors), magenta (predicted using SM samples), and yellow (predicted using FM samples). $\Delta G$ (kcal/mol) and RMSD (Å) quantify binding affinities and deviations from the experimental poses. “MD” labels protein-ligand complex MD simulations after docking.
@@ I.2 Molecular Dynamics Setup for Conformational Analysis of Protein-Ligand Complexes
I.3 Protein-ligand Blind Docking
Similarly, crystallographic water and unwanted molecules were removed from experimentally derived structures, and missing side-chains were added to generated structures via homology modeling. Receptor structures were prepared using AutoDock Tools (Morris et al., 2009), with polar hydrogens added, Kollman charges assigned, and any missing atoms repaired. For receptors within the same family, we prepared a shared ligand file, adding hydrogen atoms and assigning Gasteiger charges. A large grid box, typically 80 to 110 Å̊$\mathring{\text{A}}$ per side, was defined to cover the entire protein surface.
@@ I.3 Protein-ligand Blind Docking
Appendix J Applicability and Limitations of Deep Generative Protein Design Guidelines
  • It is hard to apply strict physical constraints to very flexible proteins in generative process. In addition, using only steric exclusions without considering environmental factors such as the lipid bilayer in membrane proteins or interfaces in large assemblies can make the designs less accurate (Winnifrith et al., 2024).
  • If a target protein’s stability or fold depends critically on bound cofactors or oligomeric assembly, designing sequences in isolation can be unreliable. According to Krishna et al. (2024), generating and screening candidate sequences without accounting for these interactions will often fail to reproduce the native structure or stability observed in the full cofactor- or multimer-associated complex.
  • For truly new folds or functions, there are no known conserved residues to guide design. In families where conserved sites are spread out or have moved over evolution, keeping those sites can block creative changes. In practice, the design of novel proteins remains a low-success “attritional” problem with success only in rare, isolated cases (Greener et al., 2018).
  • Standard force fields often do not model metal ions, sugar attachments, or membrane effects accurately. As a result, refinement steps may produce structures that differ from what actually happens in the lab.
  • For novel protein folds with no known homologs, template-based modeling or comparison is difficult. Even state-of-the-art predictors admit substantially reduced accuracy when no homologous structure exists (Jumper et al., 2021).
  • Conventional MD simulations are limited to relatively short timescales due to computational cost. For example, simulating $\sim 50,000$ atoms (a modest protein) for $\sim 1$ $\mu$s can take days on a GPU (Hollingsworth & Dror, 2018). Moreover, simulations of systems with membranes, metal centers, or covalent modifications often suffer from force-field artifacts or setup uncertainties.
  • Docking methods that do not model backbone flexibility often cannot accommodate the large backbone and side-chain rearrangements required for binding, leading to unreliable predictions (Lexa & Carlson, 2012).
Appendix K Generated Structures